Harry intends to complete a BSc degree. He has to take either Mathematics A or Physics A in his first year. If he takes Mathematics A, he can take chemistry A or Mathematics B in His second year. If he takes Physics A, he has to take Mathematics A in his second year. When Harry has to make a choice between two subjects both are likely to be chosen


1
Expert's answer
2021-04-14T11:57:07-0400

a) Draw a decision tree




b)

"P(MathA)=\frac{1}{2}"

(c)


"P(MathA\ and\ MathB)=\dfrac{1}{2}(\dfrac{1}{2})=\dfrac{1}{4}"